The present age of a man is 7x years and that of his son is (2x-1) years. When the man was (x square - 5) years old, his son was (x+1) years. Then present age of his son is-
Answer by richwmiller(17219)

7x-(2x-1)=x^2-5-(x+1)
7x-2x+1=x^2-5-x-1
5x+1=x^2-x-6
x^2-6x-7=0
(x-7)*(x+1)=0
x=7
father is 49
son is 14-1=13
check
49-5=44
8=13-5
ok